OpenStudy (anonymous):

What's the quotient to -44s^4t+77st/ 2st

OpenStudy (anonymous):

Related to your previous question \[\frac{ -44s^4 t+77st}{ 2st } \implies \frac{ -44s^4t }{ 2st }+\frac{ 77st }{ 2st }\] now you may cancel/ simplify.

OpenStudy (anonymous):

is -22s^3t+36st correct ?

OpenStudy (anonymous):